A 78.2 Kg person is standing on a scale in an elevator. If the elevator is accelerating downwards at 0.307 m/s^2 what will the reading on the scale be (in kilograms)?
1
Expert's answer
2020-10-16T11:00:58-0400

Let’s apply the Newton’s Second Law of Motion:


"\sum F_y = ma_y,""N-mg=-ma,"

here, "N" is the normal force directed upward, "mg" is the force of gravity directed downward, "a = 0.307 \ \dfrac{m}{s^2}" is the acceleration of an elevator (since the acceleration of an elevator directed downward it will be with sign minus) and "g = 9.8 \ \dfrac{m}{s^2}" is the acceleration due to gravity.

Then, we get:

"N=m(g-a) = 78.2 \ kg \cdot(9.8 \ \dfrac{m}{s^2}-0.307 \ \dfrac{m}{s^2}) = 742.35 \ N."

Finally, we can find the reading on the scale from the definition of the weight:


"m = \dfrac{N}{g} = \dfrac{742.35 \ N}{9.8 \ \dfrac{m}{s^2}} = 75.75 \ kg."

Answer:

"m = 75.75 \ kg."